13:00 UTC ITF found in SCIENCE
13:16 UTC Etalon corrections reset
13:17 UTC ITF in SCIENCE
13:45 UTC EGO truck passage near CEB
15:30 UTC ITF in COMMISSIONING
15:34 UTC inject_lsc.py started
15:36 UTC data loss (SUSP_Fb). Injections stopped
15:41 UTC injections restarted (ended 15:57 UTC)
16:00 UTC ITF in CALIBRATION (measurements with ITF unlocked)
16:03 UTC CALIBRATED_DELAY (ended 16:18 UTC)
16:21 UTC CALIBRATED_PRWI (ended 19.03 UTC)
17:14 UTC while executing READY_DF_INtoEND_LN, signal from LSC_B7_DC was a flat line and NARM_LOCK showed usermsg "Waiting for arm power estimation". Contacted ISC on call expert, NE/NI mirrors had rotated when CALI node was put in DOWN up to a point that NARM was completely misaligned. Once realigned, ITF was kept in LOCKED_ARMS_IR for a few minutes to stabilize, as instructed by the expert.
17:45 UTC started again READY_DF_INtoEND_LN. Couldn't achieve the state and LN1 in time to run CALIBRATED_DF_INtoEND_LN - Measurement of actuators response for NI,WI mirrors skipped
18:24 UTC seismic noise on B4, B7 and B8, seismon flag red, locking attempt failed. B1p_QD1 and B1p_QD2 galvo loops opened and were closed
19:12 UTC ITF in SCIENCE
19:34 UTC unlock, followed by unlocks at CARM_NULL_1F and LOCKING_CARM_NULL_3F (three times), B1p_QD1/2 Vbias rearmed
Guard Tours (UTC)
17:52 - 18:30
20:54 -
ISC
(06-10-2025 19:02 - 06-10-2025 19:40) From remote
Status: Ended
Description: NARM misalignment during calibrations, multiple unlocks
Actions undertaken: NARM showing no triggers (WARM showing no problems) after CALIBRATED_PRWI. Realigned
Status: Ended
Description: Repeated unlocks
Actions undertaken: 400 Hz oscillation spotted on mirrors